OpenStudy (anonymous):

Can someone help me determine whether the sequence converges or diverges. If it converges, give the limit. 108, -54, 27, -27/2,........ Thank you!!!!

OpenStudy (amistre64):

you would have to determine if there is some formula that can be developed to address the pattern

OpenStudy (amistre64):

the last 2 terms might give us a clue

OpenStudy (anonymous):

what you mean?

OpenStudy (anonymous):

would it diverges?

OpenStudy (amistre64):

i mean, that in order to determine what it does, it would be nice to see how the sequence is being formed. the last 2 terms that are presented give us a clue to what they are doing to form the sequence.

OpenStudy (amistre64):

what is the most reasonable explanation on how we get from: 27 to -27/2 ??

OpenStudy (anonymous):

dividing?

OpenStudy (amistre64):

yes, specifically ... dividing by -2

OpenStudy (amistre64):

does this idea hold for the other terms as well? can we get from one term to the next by dividing it by -2?

OpenStudy (anonymous):

yes, it works for the others also

OpenStudy (anonymous):

the next term would be 27/4?

OpenStudy (amistre64):

then we can assume this is a geometric sequence:\[a_n=-\frac{1}{2}a_{n-1}\] yes, 27/4 would be the next term :)

OpenStudy (anonymous):

so how do we find if it diverges or converges?

OpenStudy (amistre64):

there are a variety of methods available to us, i would prolly use a ratio test ... or, if a less stringent proof is required; 27/2^n gets very very small as n gets ver very big

OpenStudy (anonymous):

how do I use the ratio test?

OpenStudy (amistre64):

notice that out pattern is basically\[a_n=\frac{27}{2^n}\] for some starting integer value of n

OpenStudy (amistre64):

the ration a{n+1}/a{n} is the ratio test

OpenStudy (anonymous):

so the n I would plug in a number?

OpenStudy (amistre64):

\[\frac{27}{2^{n+1}}\div \frac{27}{2^n}\] \[\frac{\cancel{27}}{2^{n+1}} \frac{2^n}{\cancel{27}}\] \[\frac{2^n}{2^{n+1}}=\frac12 \] \[\frac{\cancel2^n}{\cancel2^{n}~2^1}=\frac12 \]

OpenStudy (anonymous):

A general test would be to see if your ratio has an absolute value of less than 1 (-1 <r< 1). If this is true then your series converges, that is as we get farther into the sequence the terms get closer and closer to zero.

OpenStudy (amistre64):

yes, the sequence tends to zero .... not too sure what i did there tho lol

OpenStudy (anonymous):

lol so the series would converges at 0?

OpenStudy (anonymous):

Yes. Despite the flip/flop of +/-, the terms will get smaller and smaller until, at least on your calculator, a term will be zero.

OpenStudy (anonymous):

ohh, I got it now. Thank you so much @amistre64 and @mrbarry :)

OpenStudy (anonymous):

You are very welcome. Math teachers DO want you to understand.

OpenStudy (anonymous):

how can I give you both medals? true, and I understand it now. Thank you again.

OpenStudy (amistre64):

i dont need a medal, i have enough of the useless trinkets floating about .... barry deserves to be leveled up tho

OpenStudy (anonymous):

oh, kk then and how do I give him a medal then? and thank you.